The coronavirus disease has led to many deaths globally. In Gauteng, the number of deaths per week increases by 9.2% every week. Suppose this trend does not change and 82 deaths have been recorded this week. Determine the number of deaths that were recorded 4weeks ago


1
Expert's answer
2022-03-15T16:50:41-0400

The previous period is assumed to be 100%

3rd week= "\frac{100\times 82}{109.2}= 75"

2nd week= "\frac{100\times 75}{109.2}= 69"

1st week= "\frac{100\times 69}{109.2}= 63"